OpenStudy (anonymous):

The half life for the radioactive decay of potassium-40 to argon-40 is 1.26 X 10^9 years. Suppose nuclear chemical analysis shows that there is 0.289 mmol of argon- 40for every 1.000 mmol of potassium-40in a certain sample of rock. Calculate the age of the rock. Round your answer to 2 significant digits.

OpenStudy (anonymous):

Have you tried solving it? Haw far have you got? (Use the exponential decay formula)

OpenStudy (anonymous):

I get stuck on these word problems mostly becuase it does not speak english. It all sounds like a forgien language to me.

OpenStudy (anonymous):

Is the formula A = 28(1-0.5)^42?

OpenStudy (anonymous):

The exponential decay formula is y = ae^(rt), where 'y' is the remaining amount of radioactive substance after time 't' and 'a' is the original amount of radioactive substance. Half life is given 1.26 x 10^9 years. So when t = 1.26 x 10^9, the remaining amount 'y' is half of the original amount 'a'. Hence put y = a/2 and t = 1.26 x 10^9 in the formula to get a/2 = ae^(r * 1.26 * 10^9) Can you try to solve the equation above for 'r' ?

OpenStudy (anonymous):

I have to put r on the other side of the equation in order to solve right?

OpenStudy (anonymous):

Yes, and in order to do that, you need to first remove 'a' by dividing the equation by 'a' on both sides

OpenStudy (anonymous):

Wouldn't that just make a on both sides disappear or does it put 1 in place of a?

OpenStudy (anonymous):

Yes, it would, and it is the only logical step since making 'a' disappear from the equation will have you left with only one unknown, r (e is the natural number, not a variable)

OpenStudy (anonymous):

How is e a number?

OpenStudy (anonymous):

e is the symbol for a mathematical constant just like pi

OpenStudy (anonymous):

Oh, so /2=e^(r*1.26* 10^9)

OpenStudy (anonymous):

Good! Any idea how we would solve further?

OpenStudy (anonymous):

Put 2.718281828 as the number for e?

OpenStudy (anonymous):

No you can take natural log on both sides to make e disappear

OpenStudy (anonymous):

what does natural log mean?

OpenStudy (anonymous):

Oh, do I divide both sides by e?

OpenStudy (anonymous):

no, log is a function which you can consider the opposite of 'e^()', so taking it on both sides of the equation you get log(0.5) = log(e^(r*1.26*10^9))

OpenStudy (anonymous):

value of log(0.5) is -0.693

OpenStudy (anonymous):

Use this rule now: log(e^(anything)) = (anything)*log(e), so you can simplify the right hand side to r*1.26*10^9 log(e) Now put log(e) = 1. So you get the equation -0.693 = r*1.26*10^9

OpenStudy (anonymous):

now can you solve this equation for 'r' ? This will give you the value of rate of change and hence will complete half of the question's solution

OpenStudy (anonymous):

so I again divide both sides by r wich will give us, r-0.693 =1.26*10^9

OpenStudy (anonymous):

to get the value of 'r', remove the numbers along with it. so you would divide by 1.26 * 10^9 on both sides, not r

OpenStudy (anonymous):

Oh, Ok so \[\frac{ -0.693 }{ 1.26*10^{9} }=r\]

OpenStudy (anonymous):

Good job! Need to simplify that..

OpenStudy (anonymous):

so I need to put the signintific notation in to regular numbers right?

OpenStudy (anonymous):

don't do anything with the 10^9 except bring it on the numerator by making the 9 negative,. otherwise just divide the other numbers

OpenStudy (anonymous):

I get -0.55

OpenStudy (anonymous):

Good thats what I got. Now we have the value of rate r = -0.55

OpenStudy (anonymous):

it is r= -0.55 * 10^-9

OpenStudy (anonymous):

Now we come to the scond part of the question. It says that for every 1 mol potassium - 40 you have 0.289 mol of argon. Can you guess what might be the total amount of potassium - 40 in the sample when there was no argon formed at the beginning?

OpenStudy (anonymous):

19-40 is 21

OpenStudy (anonymous):

oh, -21

OpenStudy (anonymous):

How did you get that 21? It is not given in the question

OpenStudy (anonymous):

Since 0.289 moles of argon are there it means there were as many moles of potassium at the beginning in addition to the 1 mole already present in the sample

OpenStudy (anonymous):

19 is the atomic number of potassium. I thought you were saying potassium (K) - 40.

OpenStudy (anonymous):

so 1.289

OpenStudy (anonymous):

Good!

OpenStudy (anonymous):

So we have initial amount of potassium 1.289, remaining amount of potassium 1.0 and rate -0.55 * 10^-9. Plug these in the equation Y = Ae^(rt) to get the value of t

OpenStudy (anonymous):

rt is rate I know that but is Ae the potassium or the Y?

OpenStudy (anonymous):

We are solving for Y right?

OpenStudy (anonymous):

r is the rate , t is the time required to convert 'a' moles of potassium into 'y' moles

OpenStudy (anonymous):

So Y = 1.289e^(-0.55*10^-9t) did I do that right?

OpenStudy (anonymous):

Right! and put y = 1.0 as well

OpenStudy (anonymous):

1.0 = 1.289e^(-0.55*10^-9t) and we are solving for "t" time

OpenStudy (anonymous):

yes again use the log procedure we did for the first equation to get the value of t

OpenStudy (anonymous):

log(1.0) = 1.289 log(e^(-.55*10^-9t))

OpenStudy (anonymous):

You can't just keep 1.289 out of the log just because it is a number. That does not occur with logs. Instead, to simplify things, divide by 1.289 on both sides so that you end up with a log(1/1.289) on the left hand side

OpenStudy (anonymous):

I get 461575861.6 which is 4.6*10^-8

OpenStudy (anonymous):

Excellent so you got the answer! :)

OpenStudy (anonymous):

Really! Woohoo!

OpenStudy (anonymous):

So I have one more question along these same prameters.

OpenStudy (anonymous):

Suppose the amount of a certain radioactive substance in a sample decays from 3.80 mg to 2.50 mg over a period of 49.5 years. Calculate the half life of the substance.

OpenStudy (anonymous):

Okay its the same as the previous one. First find the value of 'r'

OpenStudy (anonymous):

btw can you start a new thread please because this one is becoming too long to scroll!

OpenStudy (anonymous):

Ok, I posted it in another question.

OpenStudy (anonymous):

@Mymathhomework I'm on the new question now